40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog, Page 2 of 6  Not logged in ELOG logo
ID Date Author Type Category Subject
  217   Fri Nov 8 13:47:22 2019 Marie K.GeneralBS BRDsModel with 2 BRDs

We want to compare the model transfer function with the measurements on BS.

We built a model with two BRDs attached to the main mass. Model is provided in the attachment. Details of the computation will be given in a note later. We implemented the model in Matlab. 

Back in July (alog 177), we tuned the BRD frequencies right before mounting them on BS. We installed BRD1 and BRD3 (alog 181) with the following properties (alog 177, alog 164, alog 141):

Mode BRD1_v4 Bounce BRD1_v4 Roll BRD3_v4 Bounce BRD3_v4 Roll
Mass (g) 4.936 3.063 4.865 3.207
f (Hz) 16.693 24.26 16.66 24.32
Q 100.4 117.9 154.4 137.4
Accuracy 0.016% -0.34% -0.18% -0.097%

Plugging these properties in our new 2 BRDs model, we obtain the transfer function in Figure 1 and Figure 2. For comparison, the 1 BRD model with only BRD1(3) is shown on the same graph. The tables below show a summary of the model outputs.

Bounce F1 F2 F3 Q1 Q2 Q3
Model 2 BRDs 16.530 16.678 16.837 247 122 239
Model BRD1 16.588 16.795   202 191  
Model BRD3 16.567 16.783   267 337  
Measurement (#34) 16.55 16.64 16.75 108   172

 

Roll F1 F2 F3 Q1 Q2 Q3
Model 2 BRDs 24.051 24.289 24.583 231 128 274
Model BRD1 24.116 24.485   193 283  
Model BRD3 24.144 24.518   257 275  
Measurement (#34) 24.034   24.507 126   160

The 2 BRD model generates three poles but in the present case we observe only "2 peaks", as we have seen in the measurements (see here measurements #33 and #34). We need some larger detuning to see "3 peaks". See for example figure 3 with respectively  -0.1Hz detuning for BRD1 and +0.1Hz detuning for BRD3. The Q are ~ a factor 2 lower than expected from the model.

Attachment 1: Model3BRD_bounce_elog217.png
Model3BRD_bounce_elog217.png
Attachment 2: Model3BRD_roll_elog217.png
Model3BRD_roll_elog217.png
Attachment 3: Model3BRD_roll_detuned_elog217.png
Model3BRD_roll_detuned_elog217.png
Attachment 4: BRD_3masses_elog217.m
%% Use the results from the Model_with_2parallel_BRDs.nb
% based on the model from Norna T1600259
% MK November 2019

clc; clear all; close all;
 
% BS and damper parameters
% primary oscillator with damping
 
% BS Bounce mode
... 136 more lines ...
  216   Thu Nov 7 11:52:01 2019 Luis SanchezElectronicsSatellite Box - OSEMTest Procedure and Data Sheet

Luis

I just added a document to DCC the describe the electrical characteristics from the Aplus Satellite Amplififer E1900355. Also a test procedure was preparer to check all connections on the Aplus Satellite Amplifier E1900245.

Attachment 1: E1900355.pdf
E1900355.pdf E1900355.pdf E1900355.pdf
Attachment 2: E1900245-v2.pdf
E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf
  215   Fri Nov 1 10:40:11 2019 Craig WorleyProgress Matlab Script for Comparing Excited and Non-excited Data

Finalized revisions for version two of the Matlab script responsible for intaking B&K files and self-normalizing the different tests to allow for direct comparisons. This program's purpose is to allow for comparisons of different tests in which the excitation level is varied, or binning needs to be completed before comparisons can be made. Outputs include the normalized plot, a plot illustrating the effect of binning selection, and then a file output showing details pertaining 3dB and resonance frequency. More details can be found at T1900341, along with a demonstration video to the posted in the next week.

  214   Thu Oct 31 16:54:45 2019 Marie K.SummaryBS BRDsSummary of the BRDs results with version 4 of the blades

BRDs with the version 4 of the blades have been tested on stand-alone version and in the BS suspension during the summer. Here is a summary of our findings with the references to the corresponding elogs.

Stand alone:

Drifts measured on two BRDs before baking over 20 and 40 days (elog 186). The variation of the resonance frequencies are reported in the table below:

 

BRD1

BRD2

BRD4

 

Bounce

Roll

Bounce

Roll

Bounce

Roll

Time lapse dt [days]

19

41

5

Drift df [Hz]

0.145

0.417

0.218

0.184

0.160

0.196

Drift df [%]

0.87

1.71

0.76

1.31

0.94

0.81

Drift Rate [mHz/day]

7.63

21.96

5.32

4.49

32.00

39.20

Drift Rate [%/day]

0.09

0.04

0.03

0.02

0.18

0.16

  •  The BRD2 modes seem to stabilize after a month of steady increase (see figure 1). The rate of frequency drift is about few hundreds of ppm a day.  The drift rate of BRD1 is of the same order but the measurements were stopped before it stabilized.
  • After baking (elog 197), the BRD4 is measured over 5 days. The rate of the frequency drift dramatically increases (by a factor 4). It might be partly due to the fact that the BRD was mounted after being backed and there is a relaxation in the mount.
  • The value of the Qs is around 150 (elog 177)
  • Reference for all the measurements can be found at: T1900569

On the dummy BS suspension:

1 - Resonance frequencies

The two BRDs were installed on the dummy BS for a month. Unfortunately we are missing some of the measurements because some data got corrupted (see spreadsheet attached). Therefore the analysis is only performed over 16 days (8 days) for the bounce (roll) mode.

For each mode (bounce and roll), two peaks are observed around the resonance where we expected to resolve three peaks. The three peaks would be the main BS resonance as well as one peak per BRD. We might need to increase the scan resolution (see figure 2 and 3).
The frequency of the peaks do not match the frequency of the BRDs measured alone. The frequency of the minima in between the resonance peaks is close to the resonance frequency of the BS for respectively the bounce and the roll modes. The shift in the BS resonance frequency due to the added mass is negligible.

We observed a steady drift of the resonance frequencies over time. The frequencies are increasing by few hundreds of ppm per day, see below and figures 4 and 5.

Summary of BS Bounce frequencies

Time lapse dt [days]

16.00

16.00

Initial mistuning [%]

0.84

-0.34

Final mistuning [%]

0.64

-0.61

Drift df [Hz]

0.034

0.044

Drift df [%]

0.21

0.26

Drift Rate [mHz/day]

2.125

2.75

Drift Rate [%/day]

0.01

0.02

We are observing a change of 0.2% in the resonance frequency over the 16 days. This is above the requirements that we set at 0.1% tuning. We didn’t observe a stabilization in the drift.

Summary of BS Roll frequencies

Time lapse dt [days]

8.00

8.00

Initial mistuning [%]

1.26

-0.69

Final mistuning [%]

1.12

-0.87

Drift df [Hz]

0.034

0.045

Drift df [%]

0.14

0.18

Drift Rate [mHz/day]

4.25

5.625

Drift Rate [%/day]

0.02

0.02

We are observing a change of 0.1% in the resonance frequency over the 8 days. We didn’t observe a stabilization in the drift, so this is likely to exceed our requirements.

When remeasured stand-alone after being uninstalled from the BS suspension, we established that the BRDs frequencies drifted of about 0.5% after being for a month on the suspension (elog 203).

2 - Q factor

The resonance amplitudes, corresponding to the quality factor of the modes, fluctuate over time without a distinguishable pattern. However, it seems that for each mode the two resonances vary together, in particular for the roll modes.

The mean of the bounce mode Q is 140 (168) for the 16.66 Hz (16.69 Hz) resonance. The mean of the bounce mode Q is 115 (152) for the 24 Hz (24.5 Hz) resonance. Q are lower than expected according to the model (we expected Q~200).

Conclusion:

We observed a steady drift of the resonance frequencies of the BRDs over time, when stand alone or on the dummy BS. The frequencies are increasing by few hundreds of ppm per day. We see a stabilization in the drifts after about a month in the lab. The drift is slightly lower when the BRDs are mounted on the BS suspension compared to the stand alone BRDs in the lab. This could confirm that the excitation measurements cause some of the drift and we need to revise the method. We have no evidence that the baking process reduces the frequency drifts.

The measurements of the quality factor shows that the peaks may not be quite resolved, and that we are underestimating the Qs of the modes on the suspension. However, the value of the Qs in the stand-alone measurements is already promising.

 

Attachment 1: BRD1_drifts.png
BRD1_drifts.png
Attachment 2: Bounce_August2019_final.png
Bounce_August2019_final.png
Attachment 3: Roll_August2019_final.png
Roll_August2019_final.png
Attachment 4: BS_bounce_august_drift.png
BS_bounce_august_drift.png
Attachment 5: BS_roll_august_drift.png
BS_roll_august_drift.png
Attachment 6: August2019_monitoring.xlsx
  213   Fri Oct 18 09:42:03 2019 Luis SanchezGeneralA-plusAir Puge valve for HAM7 L1, H1 attemp number 2.

Luis, Rich, Don:

After getting some feedback from Chandra, we agree to revise the documents D1900114 and D1900116. This revision is to allocate the air purge valve on the other side of the chamber, now changing the valve from D6 Flange to D2 Flange. The documents are being modified by Don. As seen in the image, we are also adding the "HDS" on the suspension lines for better identification. The seismic db25 connectors are being allocated to D3 Flange and D3-F11 and D3-F112. The same modifications are being made to document D1900116.

Attachment 1: After_getting_some_feedback_from_Chandra.pdf
After_getting_some_feedback_from_Chandra.pdf
  212   Thu Oct 10 13:01:21 2019 Luis SanchezPlanGeneralAir Puge valve for HAM7 L1, H1

Luis:

A modification on documents D1900114 and D1900116 is on the works, and the changes will be modified by Don, as soon as he has the time. This new change is required because we need a one port for the Air Purge Valve at HAM7. This valve will help to bring the Chamber to room conditions -not under pressure-. Squeezer Tip-Tiilts that are located on D6 Flange (IC1, IC2) will change their location to D8 Flange, which is located at the top of the chamber. D8 will be a  flange with 12 dsubs with 25 pins each.

Attachment 1: Air_Purge_Valve_location_HAM7.pdf
Air_Purge_Valve_location_HAM7.pdf
  211   Wed Oct 9 16:20:01 2019 Luis SanchezElectronicsSatellite Box - OSEMInitial test from D1900217 pcb Suspension Satellite Amplifier

Luis:

Just started to test the pcb D1900217-v2, the voltage noise level appears normal and in range, see following charts. I need to take the Instrument level noise and add this to the charts.

 

Attachment 1: Noise_Test_from_LED_measured_at_R48.pdf
Noise_Test_from_LED_measured_at_R48.pdf Noise_Test_from_LED_measured_at_R48.pdf Noise_Test_from_LED_measured_at_R48.pdf
  210   Fri Oct 4 15:15:22 2019 Luis SanchezElectronicsA-plusSuspension Satellite Amplifier Chassis

Luis:

I assembled two A-Plus Suspension Satellite chassis, and also I seriliazed the pcb boards D1900217 for these D1900089 units, the numbers are S1900540 and S1900541. In the near future I will add all information on the e-traveler, for now I need to do electrical test on these units.

Attachment 1: A-plus_Suspension_Satellite_Chassis_1900089.pdf
A-plus_Suspension_Satellite_Chassis_1900089.pdf
  209   Wed Oct 2 09:53:31 2019 Luis SanchezElectronicsVariosPUM Driver Quote, Binary Output, Squeezer Wiring, and HDS

Luis:

I tried to generate a Sunstone quote for PUM Driver D070483 pcb board, but for some reason the web site was not generating the quote numbers. I asked Patrick from Sunstone and he helped us to generate the quote, now the ODB++ and gerber files are located in Sunstone server for future reference, quote number SQW-52105.

I serialized the pcb boards for Binary Output chassis, and I completed the electrical assembly of 5 units. I still need to serialized the chassis and add all information to the e-traveler. Another step that these chassis require is to do electrical test, also I need to find or to order some 1U handles, screws, etc.

I modified the Squeezer Wiring diagram by removng the VOPO and ZM Suspension elements and I updated this. I will add the removed sections to the A-plus HAM7 chamber.

Created HDS clasification table, We need to define which units will require dither and where these will be located.

Attachment 1: Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf
Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf
  208   Wed Sep 25 13:28:29 2019 Luis SanchezElectronicsVariosChassis, Test, etc

I started to assemble chassis front panel and rear panel for Suspension Satellite Amplifier unit, while pcb board is been manufactured. Suspension Satellite Amplifier rear panel was modified per Rich's comments, I need to center Local Diagnostics outputs.

I continue assembling Binary Output chassis, I found some front and rear panels and also some assembled pcbs that we will use for A+. I still need to connect the led and power on some chassis. total of 5.

Bram contacted us, because he needs PUM driver, this unit is missing a pcb document only schematics, odb and gerber files were found on DCC . I asked Sustone and Patrick mentioned to me that they can manufacture the pcb with ODB files. I updated schematic D070483 on DCC to include the driver board. 

A test was performed on Top Coil Driver with pcb D0902747, appears that circutry is susceptible to oscillations. Could not find the issue. We are thinking that the behaviour described by Dave Hoyland on received email, might be due to a load capacitance due to a long cable connection on Coil outputs, and also to op amp configuration, we saw that design use a front end unit that has high bandwidth 10MHz and unit at the back end has only 1MHz bandwidth. Rich suggested that we could use an OP-06 (slower device) in the front end (600khz bandwidth) in place of the front end op-amp.

 

Attachment 1: Preparing_received_panels_for_chassis_assembly.pdf
Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf
  207   Tue Sep 17 11:02:59 2019 Luis SanchezElectronicsA-plusChassis List and getting materials for Binary Input and Output chassis

Started to collect some items for Binary Input and Binary Output chassis, these can be see in below image. Also created document that list the preliminary chassis count for the UK electronics. This list needs to be update since the Tip-Tilts have evolve to a HDS.

Attachment 1: Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf
Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf
  206   Wed Sep 11 11:25:13 2019 Luis SanchezElectronicsA-plusAssembly Drawing A+ Suspension Satellite Amplifier

I just created an assembly drawing for the A+ Suspension Satellite Amplifier on 1U chassis.

Attachment 1: assembly_drawing_for_Sat-Amp.pdf
assembly_drawing_for_Sat-Amp.pdf
  205   Mon Sep 9 13:49:00 2019 Luis SanchezElectronicsA-plusTriple Coil Driver panels, assembly drawing and cable drawings.

After some digging on Jay Heefner old computer files, I finally found the panels for the Triple Coil drivers, I added some panel numbers and I save documents into DCC, the numbers are D1900419, and D1900418.

I also created an assembly drawing since D1001242 only shows a picture of chassis without any description, see below image.

I created a cable drawing for the cables need to connect the monitor board and to suspension satellite amplifier.

Attachment 1: D1001242_eLog.pdf
D1001242_eLog.pdf D1001242_eLog.pdf
  204   Wed Aug 28 11:20:05 2019 Andy R.ProgressBS BRDsBRD Net Drifts

Total drifts of the BRDs. BRD4 and BRD6 were baked. Baking appears to reduce v4 blade drift but increases v5 blade drift. Additionally, BRD5 has definitely stopped drifting and has settled at a max drift of around ~0.6%. 

  BRD4 Roll BRD4 Bounce BRD5 Roll BRD5 Bounce BRD6 Roll BRD6 Bounce
f Drift (Hz) 0.196 0.16 0.14 0.13 0.29 0.36
Percent Drift 0.82% 0.94% 0.56% 0.78% 1.2% 2.2%

 

Attachment 1: BRD4_Roll.png
BRD4_Roll.png
Attachment 2: BRD4_Bounce.png
BRD4_Bounce.png
Attachment 3: BRD5_Roll.png
BRD5_Roll.png
Attachment 4: BRD5_Bounce.png
BRD5_Bounce.png
Attachment 5: BRD6_Roll.png
BRD6_Roll.png
Attachment 6: BRD6_Bounce.png
BRD6_Bounce.png
  203   Wed Aug 28 11:04:00 2019 Andy R.ProgressBS BRDsBRD1 and 3 Updates

BRD1 and BRD3 Resonance peaks after being in BS Suspension for ~1 month. Note that the BRD1 roll mode sagged a lot and the peaks were did not look sharp, so drift is probably unreliable. 

  BRD1 Roll BRD1 Bounce BRD3 Roll BRD3 Bounce
Measured f(Hz) 24.39 16.59 24.45 16.73
Drift in f (Hz) from 7/31 0.133 -0.10 0.075 0.133
Percent Drift 0.55% -0.62% 0.45% 0.55%

There appears to be less total drift when left in the BS suspension when compared to the BRDs that were measured more frequently. This seems to confirm the suspicion that excitation measurements cause some of the drift. 

  202   Wed Aug 28 10:52:06 2019 Andy R.ProgressBS BRDsBRD Tuning

BRD5 and BRD6 are mounted in the BS suspension for testing and were tuned to the following resonant frequencies. Note BRD5 Roll mass was changed to: 4.23g+2.27g+long screw = 7.357g.

  BRD5 Roll BRD5 Bounce BRD6 Roll BRD6 Bounce
Measured Frequency (Hz) 24.412 16.704 24.256 16.656
Difference from Resonance (Hz) 0.072 0.014 -0.084 -0.034
Percent Difference 0.3% 0.08% -0.35% -0.2%
  201   Tue Aug 27 14:59:16 2019 Luis SanchezElectronicsA-plusSuspension Satellite Amplifier pcb test procedure

Luis:

I took some data from D1900217 and created a test procedure. Simulation results are almost in agreement with test data. Limited by SR785 when taking noise spectrum in low frequency specially between 1-10Hz.

Attachment 1: E1900245-v1_082719.pdf
E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf
  200   Mon Aug 26 14:03:43 2019 Luis SanchezGeneralA-plusFeedthroughs

Luis:

A note about suspensions feedthroughs was created to visualized the possible changes on HAM7 and HAM8, these are for electrical and vacuum  feedthroughs.

The changes include feedthroughs with 12-dB25 connectors on D3 and D5. D7 needs to be discarded since not include any form of electrical system. This changes are for H1 and L1 on HAM7 chamber.

HAM8 remains the same but need to discard D7 since this has been added to vacuum.

D8 will be for electrical.

Electrical on HAM7 and HAM8:

HAM7:

D3 12-dB25

     Filter Cavity - FC1 (D3-IC1 D3-IC2, D3-2C1, D3-2C2)

     Modified Tip- Tilt - FCR1 (D3-3C1, D3-3C2, D3-4C1, D3-4C2)

     Modified Tip- Tilt - FCR2 (D3-5C1, D3-5C2, D3-6C1, D3-6C2)

D5 12-dB25

                      SQZ  (D5-1C1, D5-1C2, D5-2C1, D5-2C2, D5-3C1, D5-3C2)

                     SPARE (D5-4C1)

       ORIGINAL TIP TILT - FCR3(D5-4C2)

       ORIGINAL TIP TILT - SQZR1(D5-5C1)

       ORIGINAL TIP TILT - SQZR2(D5-5C2)

                     SPARE (D5-6C1, D5-6C2)

D6 1-dB25 2 Fiber

                    SPARE (D6-1C1, D6-1C2)

                    SQZR 532mm (D6-2J1)

                    SQZR 1064mm (D63J1) 

HAM8:

D6 6-dB25

      Filter Cavity - FC2 (D6-1C1, D6-1C2, D6-2C1, D6-2C2)

                       SPARE (D6-3C1, D6-3C2)

D3 6-DB25

                      SPARE(D3-1C1, D3-1C2, D3-2C1, D3-2C2, D3-3C1, D3-3C2)

Vacuum Elements on HAM7 and HAM8:

D7 for ION PUMP,

C1 for TURBO PUMP and PRESSURE GAUGE, this will be implemented via "T" with a nipple for the pressure gauge.

D8 ELECTRICAL

RGA will be located on door at -x axis.

 

 

 

Attachment 1: Feedthrough_arrangement.pdf
Feedthrough_arrangement.pdf Feedthrough_arrangement.pdf
  199   Wed Aug 21 17:02:21 2019 Luis SanchezElectronicsA-plusSuspension Satellite Amplifier pcb evaluation

Luis:

Recieved the pcb from screaming circuits and I found out that they installed the incorrect device in OP2177 places, 12 units were affected, they installed a SPI Flash Memory Adesto 1749 25DF081A. Screaming circuits is reviewing to see what have failed in the process, but appears that Digi-Key made a mistake since the bag containing the chips are label correctly but the parts inside are the memory ic.

I added the missing 20uF capacitors on all 4 channels.

Started to test and identified some issues with this board, fixing the errors and verifiying DC voltages.

Attachment 1: Suspension_Satellite_Amplifier_PCB_evaluation.pdf
Suspension_Satellite_Amplifier_PCB_evaluation.pdf Suspension_Satellite_Amplifier_PCB_evaluation.pdf Suspension_Satellite_Amplifier_PCB_evaluation.pdf Suspension_Satellite_Amplifier_PCB_evaluation.pdf
  198   Wed Aug 21 11:19:48 2019 Andy R.ProgressBS BRDsBRD6 Masses

[Agueda, Andy]

BRD6_v5 BRD6_v5 BRD4_v4 BRD4_v4
Roll Mode Mass Bounce Mode Mass Roll Mode Mass Bounce Mode Mass
4.17g+2*1.02g+long screw = 7.310g 4.23g+2.38g+2.27+long screw = 10.063 2.70g+screw = 3.188g 2.74g+1.57g+screw=4.915g

 

  197   Mon Aug 19 16:21:08 2019 Marie K.ProgressBS BRDsBaking the blades

[Calum, Agueda, Marie]

We are looking into solutions to the sagging of the blades (v4 and v5) over time as measured by Andy. One idea is to bake the blades as it would be done to prepare them for the installation at the sites. Today we started the oven and set it to 120 deg C. We will bake the blades for 48 hours as described here for example.

  • 08/19/19 1:00 pm: 48 degC. Two v4 blades and two v5 blades were added in the oven.
  • 08/19/19 2:00 pm: 106 degC
  • 08/19/19 4:00 pm: 106 degC
  • 08/20/19 7:00 am: 106 degC. The oven temperature is low but we will leave it this way for now.
  • 08/21/19 9:00 am: 106 degC. We turned off the oven and took the blades out.

The blades will now be used to build new BRDs and check if the drifing is reduced by the bake.

 

  196   Mon Aug 19 15:57:52 2019 Marie K. ProgressBS BRDsResolving the resonance peaks?

We are currently tracking the evolution of the resonances over time. As Norna pointed out, I am taking measurements of the resonances with a coarser resolution than she did. It means that the peaks may not be quite resolved, and that we are underestimating the Qs of the modes.

On Friday, I measured the lower peak of the bounce mode with a higher resolution. I did this measurement just after the usual measurement of the two bounce peaks (see figure 1). The results from the two measurements are similar in frequency and Q. We will repeat with other resonances before drawing definitive conclusions.

Attachment 1: Meas_54_55.png
Meas_54_55.png
Attachment 2: Meas_54_55_zoom.png
Meas_54_55_zoom.png
  195   Mon Aug 19 14:34:08 2019 Agueda B.SummaryIR cameraMillimeter/Pixel Calculations for the Point Absorbers

Below are my calculations to find the millimeter/pixel (mm/pxl) within two images taken by the IR camera: one of the first point absorber and one of the second point absorber that were previously on the test mass. Attached are the pictures used to find the mm/pxl results. 

 

PT Absorber 1 

*The pythagorean theorem was used only for the first point absorber's calculations, as the pixel ruler on Preview would not work because the mark is angled. Changing the angle of the picture would change the picture size, thus producing an inaccurate mm/pxl result*

(Marker as a Whole)

- 9.5 mm/ 251 pxls wide = 0.04 mm/pxl

- 6.0 mm/ 153 pxls tall =  0.04 mm/pxl

(Inner Circle of Marker)

- 2.5 mm/ 73 pxls wide =  0.03 mm/pxl

- 1.9 mm/ 51 pxls tall =  0.03 mm/pxl

(Marker Outline)

- 3.0 mm/  93 pxls thick (left) =  0.03 mm/pxl

- 3.1 mm / 97 pxls thick (right) =  0.03 mm/pxl

- 1.9 mm/  52 pxls thick (top) =  0.04 mm/pxl

- 1.5 mm/  45 pxls thick (bottom) =  0.03 mm/pxl

 

PT Absorber 2

(Marker as a Whole)

- 4.5 mm/ 136 pxls wide =  0.03mm/pxl

- 5.8 mm/ 170 pxls tall =  0.03 mm/pxl

(Inner Circle of Marker)

- 1.5 mm/ 55 pxls wide =  0.03 mm/pxl

- 2.0 mm/ 62 pxls tall =  0.03 mm/pxl

(Marker Outline)

- 1.6 mm/ 52 pxls thick (left) =  0.03 mm/pxl

- 1.0 mm/ 45 pxls (right) = 0.22 mm/pxl

- 1.9 mm/  60 pxls (bottom) = 0.03 mm/pxl

- 1.5mm/ 58 pxls (top) =  0.03 mm/pxl

Attachment 1: First_Point_Absorber_(August_6).tiff
Attachment 2: Second_Point_Absorber_(August_6).tiff
  193   Mon Aug 19 12:22:07 2019 Andy R.ProgressBS BRDsDrift Updates

The v5 of the BRD continues to drift, the following is the total net drift from 8/1/19 to 8/19/19:

BRD5_v5 Roll Mode Bounce Mode
Net Drift 0.128 Hz 0.128 Hz
Percent Drift 0.512% 0.777%

 

Attachment 1: roll.png
roll.png
Attachment 2: bounce.png
bounce.png
  192   Mon Aug 19 12:10:01 2019 Agueda B.SummaryIR cameraRemeasuring the Beam Size and Finding the Second Point Absorber

[Marie, Andy, Agueda]

Using a beam analyzer, we found that the laser beam had a mean width (diameter) of 5438.17 micro meters. Below are the images produced by the beam analyzer. 

We then proceeded to record the second point absorber on the test mass with the IR camera. Although we increased the beam's percent power every minute, with the maximum being at 95%, the laser did not show up on the recording.

Attachment 1: meas1_20190819.png
meas1_20190819.png
Attachment 2: meas2_20190819.png
meas2_20190819.png
Attachment 3: meas3_20190819.png
meas3_20190819.png
  191   Thu Aug 15 08:19:04 2019 Luis SanchezElectronicsGeneralVCO voltage source or low pass filter?

Luis:

A fixture to energize the 800MHz VCO was assembled, this unit was implemented with a 9v battery,a mechanical potentiometer and a low pass filter. After testing this fixture we realized that the mechanical potentiometer is inducing extra noise at low frequencies. A new version of the device was assembled, only the filter was kept, now we will proceed to test the VCO with the Valon 3010, to lower frequency to 80MHz. To power the modified box we use a Martell IVC-222HPII power source.

Attachment 1: VCO_tune_voltage_source_assembly.pdf
VCO_tune_voltage_source_assembly.pdf VCO_tune_voltage_source_assembly.pdf
Attachment 2: noise_batt.pdf
noise_batt.pdf
Attachment 3: noise_martell.pdf
noise_martell.pdf
  190   Tue Aug 13 10:08:53 2019 Marie K.How ToIR cameraIncident power calibration

[Agueda, Andy, Marie]

Using a powermeter, we measured the incident power at the test mass as a function of the generated laser power. This work has already been mentioned in elog 188, but we didn't extract the calibration at that time. The power meter can measure values up to 10W at 1064nm.

  • We found that the incident power is 90% of the generated power as displayed on the laser controller (see text file attached).
  • For practicity reasons, we derived the incident power on the test mass as a function of the requested power in %. The calibration is the following: P_testmass = 0.5*P_% -4.4. This is summarized in the graph attached.
Attachment 1: Power_calibration.png
Power_calibration.png
Attachment 2: 20190809_laser_power.txt
Laser power (in W):
Ouptut	Input	Percent Power
0.335	0.3	9.5%
0.57	0.6	10%
1.57	1.7	12%
3.09	3.4	15%




... 34 more lines ...
  189   Fri Aug 9 17:46:04 2019 Marie K.ProgressIR cameraRemoving the marker tracks from the test mass coating

[Agueda, Andy, Liyuan, Marie]

To be able to observe point absorbers, we have to remove the marker prints. On Tuesday Liyuan applied first contact to the test mass surface. After 2-3 hours of drying, we were able to peel-off the first contact. It completely removed the marks from the test mass surface. Pictures can be found at: https://photos.app.goo.gl/sjfkWssdm1mEKh3FA

Beforehand, we took reference of the center of the marks, where the point absorbers should be:

  • Point absorber 1 reference {-28.8440, 89.9750}
  • Point absorber 2 reference {7.4980, 85.1125}
  188   Fri Aug 9 16:31:13 2019 Marie K.ProgressIR cameraFirst observations with the IR camera

[Agueda, Andy, Marie]

We looked at the two zones where we expect to see point absorbers. These zones had been previously identified by GariLynn and Liyuan, and the camera vertical position adjusted to see clearly the test mass surface (see elog 189 and elog 182).

We started to develop a protocol to track the point absorbers. The idea is to look for changes in temperature related to the increase of input power. We gradually increased the input power and looked at the temperature image difference from the initial "cold" state. The exact timing of our measurements today is attached in the text file. The camera was set to take an average of of the temperature at 4Hz. We recorded a picture every second.

First zone (coordinates {-28.8440, 89.9750}): we started to see a clear point in the middle of the zone at 35%-40% input power (to be converted in W). When turning off the laser, the point disappeared rapidly (timescale < second). The temperature of the point increased with the power of the laser. We could see more than 1 deg increase on one pixel at max. Further characterization needed.

Second zone (coordinates {7.4980, 85.1125}): we couldn't see any point even at the maximum input power (70%). An offline analysis might help to dicern features.

Unfortunately, the recordings of our experiment are not saved in the format we were expected. Only the temperature profiles are saved, not the entire image. There is not much we can do with this, and we will have to retake the measurements on Monday. Yesterday, we took a picture and a video file of the first zone under 50% of input power with a shorter frame rate (20Hz?). The picture is quite noisy, but we can see a increase of temperature in the center of the picture.

As a side note, we believe that we can see the beam dump gradually heating up on the left side of the camera. Also converting a RGB picture into a grayscale picture makes use loose the temperature calibration, so we should find another solution.

To do: retake the same measurement, find the zones with temperature gradients and extract the temperature over time.

Power calibration: We used a power-meter (thanks Liyuan!) to know the ratio of input power read on the laser controller display versus laser power on the test mass. The calibration is attached in the text file.

Attachment 1: differerence_test.png
differerence_test.png
  187   Thu Aug 8 09:35:17 2019 Marie K.ProgressBS BRDsMeasurements of the BS mode with BRDs

We are still measuring the BS modes with BRD1 and BRD3 (v4). We are observing changes in the resonances over time, which are probably consistent with the drifts Andy is reporting on the stand alone BRDs (see elog 186).

Attachment 1: Bounce_August2019.png
Bounce_August2019.png
Attachment 2: Roll_August2019.png
Roll_August2019.png
  186   Wed Aug 7 10:22:53 2019 Andy R.ProgressBS BRDsUpdated BRD Drifts

Net drifts for BRDs:

  BRD1_v4 Bounce BRD1_v4 Roll BRD2_v4 Bounce BRD2_v4 Roll BRD5_v5 Bounce BRD5_v5 Roll
Total Drift (Hz) 0.145 .0417 0.202 0.152 0.072 0.068
Percent Drift 1.211% 1.712% 0.872% 0.627% 0.437% 0.272%

 

Attachment 1: bounce.png
bounce.png
Attachment 2: roll.png
roll.png
Attachment 3: v5_roll.png
v5_roll.png
Attachment 4: v5_bounce.png
v5_bounce.png
  185   Tue Aug 6 13:28:38 2019 Luis SanchezElectronicsValon 5009Synthesizer Valon 5009

Luis:

Valon 5009 Unit 2.

This unit failed after power up, it is no locking and is no putting the correct output frequency, I  do not know when the unit was damaged. This device was used on the cryogenic lab in bidge building.

Some symptoms: when the unit is locked usually the lock button change color from red to green indicating that device is lock. Well this unit at power up sometimes pulls around 200mA with 6v or slightly less; after set the frequency to 80MHz the gui display a comment of 80MHz but the lock button never change color from red to green. One thing I noticed is that when pressing the rf output and the synth power buttons  and leaving both in off state, and then enable the synth power button unit change its state on gui by showing the lock button green. Another is tha the output signal sometimes does not correlate with the frequency setting or attenuation setting on gui and signal output. 

Unit was shipped to Valon Technologies at Redwood CA, on August 6th 2019. The unit is no longer under warranty but Stuart Rumley offer to check the unit to see if he can find the problem.

Attachment 1: Valon_5009_unit_2.pdf
Valon_5009_unit_2.pdf
  184   Mon Aug 5 14:32:32 2019 Luis SanchezElectronicsPDHTransfer Fuunction Simulation Results with new topoly

Luis:

A few weeks ago, I received a new topology for Transfer Function (Johannes modification) circuit; this board still has 5 op-amp stages that will be activated in a different way. Still 3 switches but the activation is as follow: SW1 activate OP AMP1, SW2 activate OP AMP2 and OP AMP3, SW3 activate OP AMP4 and OP AMP5. I ran some simulations and the results are below. I need to claryify some points with Johannes since some simulated results are different, appears that switch activation was run in a different way for some plots, but the final configuration appears ok.

Attachment 1: Test_Simulation_Results_from_modified_D1700471_circuit.pdf
Test_Simulation_Results_from_modified_D1700471_circuit.pdf Test_Simulation_Results_from_modified_D1700471_circuit.pdf Test_Simulation_Results_from_modified_D1700471_circuit.pdf
  183   Mon Aug 5 13:56:13 2019 Agueda B.ComputingBS BRDsMetal Density Comparison for BRDs

The images below are computations to try to find the length of a BRD in different metals that would have the same weight and top square dimentions as the current Copper BRDs but with a shorter length than the current ones (0.865 cm). 

(These calculations were made without taking the hole in the center of the BRD into account)

Here are the lengths that BRDs of various metals would need to be in order to meet the above requirements:

  • Steel
    • Grades: 302, 304, 304L, 321
      • 0.975 cm
    • Grades: 310s, 316, 316L, 347
      • 0.970 cm
    • Grades: 409, 430, 434
      • 1.005 cm
    • Grades: 405, 410, 420
      • 0.998 cm
  • Aluminum 
    • 2.764 cm
  • Lead 
    • 0.682 cm
  • Tungsten
    • 0.400 cm
  • Bismuth
    • 0.776 cm
Attachment 1: BRD_Sketch.jpg
BRD_Sketch.jpg
Attachment 2: Steel_Length_Calculations.png
Steel_Length_Calculations.png
Attachment 3: Aluminum_Length_Calculations.jpg
Aluminum_Length_Calculations.jpg
Attachment 4: Lead_Length_Calculations.jpg
Lead_Length_Calculations.jpg
Attachment 5: Tungsten_Length_Calculations.jpg
Tungsten_Length_Calculations.jpg
  182   Mon Aug 5 13:06:37 2019 Agueda B.ProgressIR cameraSetting the camera realtive to the test mass

We re-installed the camera, mirrors, and beam bump to their previous positions in the RTS.

Adjustments had to be made to the angle of the beam bump with the laser at 5% before the camera was turned on. Once the wires were attached to the camera and the laptop, the laser beam power was increased to 9.5% and the camera was turned on, producing the images below.

The translation plate and camera were adjusted and moved to produce clearer images of the test mass. We were able to clearly see the marker tracks on the surface of the test mass, with and without the laser on. This set our reference for the translation stage position (0.833). For the last two images, 6 and 7, we drastically reduced the range of the color display of the image to make the image more distinct (this is mislabeled as "change of exposure" but it is not actually changing the frame rate). The emissivity of the marker tracks is different from the test mass coating, so our pictures are dominated by the tracks radiation when the laser is on. We will remove the marks with first contact asap.

Once the images were saved on the computer, we disconnected the computer from the camera and left it, along with the mirrors and beam bump, on the breadbord in its position in the RTS for more point absorber study tomorrow. 

Attachment 1: First_Point_Absorber_August_5_2019.tiff
Attachment 2: First_Point_Absorber_2_August_5_2019.tiff
Attachment 3: First_Point_Absorber_3_August_5_2019.tiff
Attachment 4: First_Point_Absorber_4_(background)_August_5_2019.tiff
Attachment 5: First_Point_Absorber_5_(minus_3_cm_y)_August_5_2019.tiff
Attachment 6: First_Point_Absorber_6_(change_of_exposure)_August_5_2019.tiff
Attachment 7: First_Point_Absorber_7_(change_of_exposure)_August_5_2019.tiff
Attachment 8: First_Point_Absorber_8_(reference_for_center)_August_6_2019.tiff
Attachment 9: Second_Point_Absorber_2_(referencefor_center)_August_6_2019.tiff
  181   Sun Aug 4 16:36:07 2019 Marie K.SummaryBS BRDsSummary of BS measurements after retuning

Andy retuned BRD1_v4 and BRD3_v4 after they had been left alone for a week in the lab. The tuning is described here: alog 177. It is better than 0.4% for all modes of the BRDs. I reinstalled them on BS right afterwards and took measurements of the bounce and roll modes over the following days:

We can see (figure 1) that the resonances are drifting. The BS bounce mode is 16.69 Hz without BRDs, so it is probably B2 here with the added mass.

Bounce

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) BRD tuning Measured range  Nb points Drive [V] Filename Comments
Measurement B12 16.550 108 16.747 172 - 16 - 17 600 1 srs0034.txt BRD1_v4 & BRD3_v4
Measurement B13 16.555 153 16.761 232 - 16 - 17 600 1 srs0041.txt BRD1_v4 & BRD3_v4
Measurement B14 16.560 139 16.754 227 - 16.4 - 17 600 1 srs0042.txt BRD1_v4 & BRD3_v4

 

The roll mode also experiences drifts, but to a lesser extent. The BS roll mode is 24.34 Hz without the BRDs. It is probably R2 here with the added mass.

Roll

R1 Freq. [Hz]

Q1 (from 3dB width) R2 Freq. [Hz] Q2 (from 3dB width) BRD tuning Measured range  Nb points Drive [V] Filename Comments
Measurement R13 24.034 126 24.507 160 - 24 - 25 600 1 srs0033.txt BRD1_v4 & BRD3_v4
Measurement R14 24.049 138 24.526 187 - 23.5 - 25.5 500 1 srs0039.txt BRD1_v4 & BRD3_v4
Measurement R15 24.051 139 24.528 195 - 23.5 - 25.5 600 1 srs0040.txt BRD1_v4 & BRD3_v4

 

Unfortunately, some of the measurements are not readable. I will retry the conversion from .78D to .txt. Files I have to mention that saving the data from the spectrum analyzer is unreliable at times, I already lost few hours of measurements because the files were not saved correctly into .78D in the first place (.78D files were empty).

Attachment 1: B12_B14.png
B12_B14.png
Attachment 2: R13_R15.png
R13_R15.png
  180   Sun Aug 4 16:24:54 2019 Marie K.How ToIR cameraConnection to the IR camera

Agueda and Andy worked out how to connect the IR Optris camera to a laptop.

Quick Start Infrared Camera PI 640:

  1. Start Software PIX
  2. Plug in the PIF cable to the Right M4 thread hole of the camera
  3. Plug in the USB cable to the Left M4 thread hole of the camera and then to the computer
  4. An image should appear on the software

Quick Shutdown:

  1. Remove the PIF cable
  2. Unplug the USB cable from the computer and then from the camera
  3. Quit Software PIX
  179   Fri Aug 2 16:44:07 2019 Andy R.ProgressIR cameraPixel Size

Image of mm ruler using the IR camera. The pixel size is \bigg(\frac{26\text{ pixels}}{1\text{ mm}}\bigg)^{-1} = 0.3846 \frac{\text{mm}}{\text{pixel}}. Edit Marie: I think there is a factor 10 to correct here (38 um/pxl).

  178   Fri Aug 2 14:58:50 2019 Marie K.ConfigurationIR cameraInstallation on RTS bench

[Agueda, Andy, Marie]

We prepared the IR camera set up in the lab B137 for the point absorber study. To integrate our setup inside the RTS and make it easier to remove when needed, we use a 6"x6" breadbord provided by Liyuan.

The tentative setup is the following:

  • The IR camera (Optris 640PI) is on a translation stage on the breaboard for easier height adjustment. The camera is placed 4" above the test mass.
  • There are two additional steering mirrors to direct the 1064 nm laser beam on the test mass below the camera. The first steering mirror M1 is on the main bench. The second sterring mirror M2 is on our breadboard.
  • A beam dump is placed in the reflection path of the 1064 nm laser beam. Due to a lack of space on the breadboard, the beam dump is installed on a extra base screwed on the back of the bench.

We turned on the laser at 0.6 W and checked that the beam goes up to the camera field of view.

The camera and the beam dump have been removed from the RTS bench for the weekend, because we would like to have Liyuan approval before.

Attachment 1: Setup_overview.png
Setup_overview.png
Attachment 2: Setup_detail.jpg
Setup_detail.jpg
  177   Wed Jul 31 11:57:19 2019 Andy R.ProgressBS BRDsBRD1 and BRD3 Tuning

Retuning of BRD1_v4 and BRD3_v4 Bounce and Roll modes. Given results are averages over 4 measurements and accuracy is from desired frequencies of 16.69 Hz and 24.34 Hz for the Bounce and Roll modes, respectively. 

Mode BRD1_v4 Bounce BRD1_v4 Roll BRD3_v4 Bounce BRD3_v4 Roll
f (Hz) 16.693 24.26 16.66 24.32
Q 100.4 117.9 154.4 137.4
Accuracy 0.016% -0.34% -0.18% -0.097%

 

  176   Tue Jul 30 21:09:20 2019 Marie K.ProgressBS BRDsBounce mode of the beam splitter

Measuring the bounce mode of the beam splitter alone, the frequency of resonance is 16.69 Hz with a Q of ~3000 (as expected).

Bounce

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) BRD tuning Measured range  Nb points Drive [V] Filename Comments
Measurement B9 16.6897 2868 - - - 16.675 - 16.710 600 0.1 srs0026.txt no BRD
Measurement B10 16.6896 3104 - - - 16.675 - 16.710 600 0.1 srs0027.txt no BRD
Measurement B11 16.6898 2807 - - - 16.675 - 16.710 600 0.1 srs0028.txt no BRD
Attachment 1: B9toB11.png
B9toB11.png
  175   Tue Jul 30 13:36:32 2019 Andy R.ProgressBS BRDsk Damping Coefficient

The damping coefficient for the v4 blade is  1.48 \text{ N/m }\pm 0.265 \text{ N/m}  from 8 different mass measurements. The damping coefficient of the v5 blade is 5.19 \text{ N/m }\pm 2.45 \text{ N/m} from 8 different mass measurements. 

  174   Mon Jul 29 22:46:11 2019 Andy R.ProgressBS BRDsMeasuring BRD Drifts

Drifts of the BRD1_v4 and BRD2_v4 over the span of ~3 weeks. Will update over the next 2-3 days to see if drift is continuing or stabilized. All data is in Hz and averaged over 4 measurements. 

Day BRD1_v4 Roll BRD2_v4 Roll BRD1_v4 Bounce BRD2_v4 Bounce
7/11   24.236   16.678
7/12 24.372   16.636  
7/18 24.6955 24.296 26.758 16.76
7/29 24.76975 24.368 16.7735 16.844

 

Net Drift

  BRD1_v4 Roll BRD2_v4 Roll BRD1_v4 Bounce BRD2_v4 Bounce
Total Change (Hz) 0.39775 0.132 0.1375 0.166
Percentage Change 1.632% 0.545% 0.827% 0.995%

 

  173   Mon Jul 29 09:55:08 2019 Marie K.ProgressBS BRDsBS resonances

Adding a summary figure with the data extracted from the files. The Q on the figure are slightly different from the one reported in the table, because they are now computed from extrapolating the data in Matlab (should be more precise than reading the values on the spectrum analyzer).

Quote:

Starting to measure the BS resonances more finely.

Bounce Freq [Hz] Q (from 3dB width) Measured range Nb points Drive [V] Filename Comments
Measurement B1 16.6937 +/- 0.0008 909 16.5 - 17.0 600 0.1   no BRD
Measurement B2 16.6922 +/- 0.0001 2200 16.675-16.710 600 0.1   no BRD

The measured Q is low for the first measurement, but might be due to the low resolution. Measurement B2 is closer to what is expected from previous measurements (see Norna's document T1700176-v7, page 4): bounce mode frequency = 16.70 Hz, Q = 3400. One or two more measurements would be useful to confirm the Q.

 

Roll Freq. [Hz] Q (from 3dB width) Measured range  Nb points Drive [V] Filename Comments
Measurement R1 24.341 +/- 0.003 - 24 - 25 600     no BRD
Measurement R2 24.33477 +/- 0.00005 2960 24.320 - 24.350 600 0.01 SRS001.78D no BRD
Measurement R3 24.33722 +/- 0.00003 3320 24.330 - 24.345 600 0.01 SRS003.78D no BRD
Measurement R4 24.33601 +/ 0.00003 3294 24.330 - 24.345 600 0.05 SRS004.78D? no BRD
Measurement R5 24.33441 +/- 0.00003 2500 24.330 - 24.345 600 0.03 SRS005.78D no BRD
Measurement R6 24.33619 +/- 0.00003 2884 24.330 - 24.345 600 0.01 SRS006.78D no BRD

To be compared to Norna's results: Roll mode frequency = 24.34 Hz, Q = 3000.

 

Attachment 1: BS_roll_noBRD.png
BS_roll_noBRD.png
  172   Thu Jul 25 08:52:55 2019 Luis SanchezElectronicsPDHTransfer Function Board DC Test - MIT Nick's boards

Luis:

A test was conducted to verified the behaviour observed by Nick at MIT. "High current on TF board"

From inspection and dc test conducted, the unit "yes" saturate over time, and make this saturation worse when all integrators are enable at the same time. Not high current was observed during test, only the incremental current due to the saturation proper of the op amp integrator action.

Attachment 1: MIT_student_intern_Nick_P__TF_board.pdf
MIT_student_intern_Nick_P__TF_board.pdf MIT_student_intern_Nick_P__TF_board.pdf MIT_student_intern_Nick_P__TF_board.pdf MIT_student_intern_Nick_P__TF_board.pdf
  171   Wed Jul 24 10:53:08 2019 Marie K.How ToBS BRDsRead data from spectrum analyzer saved on floppy disk

The current procedure to read the data is:

  1. Transfer the data from the floppy disk to a Windows computer. For now I have to use the floppy disk reader on my Mac, save the .78D files on a USB key. Then save the .78D from the USB key on my Windows computer.
  2. Convert the files from .78D in .txt files by:
    1. on a windows computer, download the SR785 software from Stanford Research Systems (currently (July 2019) https://www.thinksrs.com/downloads/soft.html)
    2. move the .78D files in the SR785 software directory
    3. open a terminal and go to the SR785 directory
    4. convert the files from the terminal using the command : >> srtrans.exe SRS001.78D srs001.txt
  3. Then the .txt files can be read with Matlab. Their header contains the spectrum analyzer configuration for the measurement.
  170   Wed Jul 24 10:13:38 2019 Marie K. SummaryBS BRDsSummary of tests with BRDs v4

I haven't posted the BS transfer functions results since we started to test the BS with 2 BRDs attached. I can now read the data that has been saved in .78D from the spectrum analyzer, which makes data analysis much easier! (see elog 171).

  • The first set of tests (tests srs0014.txt to srs0018.txt) were done with BRD1_v4 and BRD2_v4 on the beam splitter.
  • Then, I noticed the blade of BRD2_v4 had been torqued to much. BRD2_v4 was replaced by BRD3_v4. The tests are srs0019.txt to srs0021.txt.
  • Finally I removed both BRDs from the BS to remeasure their tuning. I took this opportunity to remeasure BS modes undamped because I didn't have global scans saved in the .78D format. I'm still missing precise scans for the bounce mode (also because not saved in .78D format originally).

The results are summarized in the tables below and in the figures attached.

Bounce

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) BRD tuning Measured range  Nb points Drive [V] Filename Comments
Measurement B6 16.733 325 - -   16 - 17 600 1 srs0015.txt BRD1_v4 + BRD2_v4 on
Measurement B7 16.742 46 - - BRD3 tuned 16.3 - 17.2 600 1 srs0020.txt BRD1_v4 + BRD3_v4 on
Measurement B8 16.749 289 - - drift from B7? 16 - 17 600 1 srs0021.txt BRD1_v4 + BRD3_v4 on

Is the low Q of measurement B7 due to a very good tuning? See elog 164, BRD3 was tuned to 0.01%. The measurement B7 was taken soon after tuning, and the BRDs might have drifted for the folllowing measurement B8. However, we don't see this good performance in the overall scan that was taken just after installing BRD3 on BS (see figure 3).

To do: retake transfer functions of the bounce mode region with and without the dampers to confirm the Qs. Understand the difference in the measurements (drifts?). A tentative summary is presented in figure 1, but the undamped mode is missing from the figure.

 

The results are so far are cleaner for the roll mode (see figure 2).

Roll

R1 Freq. [Hz]

Q1 (from 3dB width) R2 Freq. [Hz] Q2 (from 3dB width) BRD tuning Measured range  Nb points Drive [V] Filename Comments
Measurement R11 24.116 182 24.419 161   24 - 24.6 600 0.1 srs0016.txt BRD1_v4 + BRD2_v4 on
Measurement R12 24.110 174 24.414 167   24 - 25 600 0.1 srs0018.txt BRD1_v4 + BRD2_v4 on

We can see that adding a second BRD helped to reduce the Q of the second resonance from ~250 to ~170. From elog , we know that the undamped Q of BS is around 3000, and that adding the first BRD split the resonance in two, with Q1 below 200 and Q2 around 250. A third resonance/feature can be seen on the glocal scans, but it is outside the frequncy range of our current measurements.

To do: Take measurements over a larger range (on the high frequency side) to check the third resonance. Add a comparison with the model to check if the measured behavior is expected.

 

Attachment 1: BS_bounce_BRD_July2019.png
BS_bounce_BRD_July2019.png
Attachment 2: BS_roll_BRD_July2019.png
BS_roll_BRD_July2019.png
Attachment 3: BS_summary_July2019.png
BS_summary_July2019.png
  169   Fri Jul 19 11:28:11 2019 Luis SanchezElectronicsA-plusSuspension Feedthroughs

Luis

Started to count all needed feedthroughs for A+ suspensions and I created the next excel sheet for reference. With this new list I will verify the Bill of Materials for all suspension structures needed for A+, and probably will need to modify the already created BOM lists and altium diagrams.

Attachment 1: FEEDTHROUGHS_LIST.pdf
FEEDTHROUGHS_LIST.pdf FEEDTHROUGHS_LIST.pdf
  Draft   Thu Jul 18 20:41:23 2019 Andy R.ProgressBS BRDsBRD1, 2, 3

These are the BRD resonant frequencies as of the afternoon of 7/18: (averaged over 4 measurements)

  BRD1_v4 Bounce BRD1_v4 Roll BRD2_v4 Bounce BRD2_v4 Roll BRD3_v4 Bounce BRD3_v4 Roll
f (Hz) 16.758 24.70 16.76 24.296 16.636 24.328
Q 146.66 149.9 139.17 124.49 153.73 n/a
Measurement # 47-50 51-54 137-140 141-144    

Will measure again in about 10 days to see long term drift.

  166   Thu Jul 18 15:24:38 2019 Luis SanchezElectronicsA-plusPCB A+ Suspension Satellite Amplifier

Luis

I finished the pcb design of the 1U A+ Suspension Satellite Amplifier circutry, I placed the order in screaming circuits and in around 2-3weeks the electrical testing will start. The attached is a 3D image generaded from Altium. I am also adding the Draftsman file also created in Altium. I will add all documents to DCC D1900217.

 

 

Attachment 1: 3D_image.pdf
3D_image.pdf 3D_image.pdf
Attachment 2: SatBox_DRAFTSMAN_D1900217.pdf
SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf SatBox_DRAFTSMAN_D1900217.pdf
ELOG V3.1.3-